OpenStudy (anonymous):

x^2-6x+8=0 x^2-4x-5=0 x^2+8x+15=0 -x^2-2x+3=0 x^2+4x+8=0 -x^2+2x=12 x^2-4x+2=0 x^2-6x+9=0

OpenStudy (anonymous):

I have to graph the answer @thomaster

thomaster (thomaster):

Do you need to solve for x?

OpenStudy (anonymous):

yes @thomaster

thomaster (thomaster):

okay the first one \(x^2-6x+8=0\) You either need to get binomials in parentheses or use the quadratic formula. The general form of these quadratic equations is \(ax^2+bx+c=0\) To get the stuff in parentheses you need to find 2 numbers that are bx when added and c when multiplied. For the example we need to find 2 numbers who are -6 when added to eachother and 8 when multiplied. It's just a little bit of trial and error. I can see that it has to do something with 4 and 2 because 4*2 = 8 and 4+2 = 6. the only problem is that its -6 and not 6. So you play a little bit with the numbers: -4*2 = -8 and -4+2 = 2 so this is not correct 4*-2 = -8 (you can see it's incorrect already here because it needs to be 8) -4*-2=8 and -4+-2=-6 (+- is the same as -) Now you have the numbers and put them in parentheses like this: \(\Large(x-4)(x-2)=0\) you know this is correct by distributing the binomials again, you'd get \(x^2-6x+8=0\) to continue: \((x-4)(x-2)=0\) \((x-4)=0~~and~~(x-2)=0\) \(x-4=0~~and~~x-2=0\) first one you add 4 on both sides and 2 at the second \(x=4~~and~~x=2\) \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)